The age of the father is 5 times the age of his son. After 18 years, the age of the father will be twice his sons age. Find their present ages.
Answer by Maths68(1474)

Let
Present age of Father = f
Present age of Son = s
The age of the father is 5 times the age of his son.
f=5s.........(1)
After 18 years,
Age of Father = f+18
Age of Son = s+18
Then, the age of the father will be twice his sons age.
f+18=2(s+18)
f+18=2s+36
f=2s+36-18
f=2s+18............(1)
Substitute the value of f in (2) from (1)
5s=2s+18
5s-2s=18
3s=18
3s/3=18/3
s=6
Substitute the value of s in (1)
f=5s
f=5(6)
f=30

Present age of Father = f = 30 years old
Present age of Son = s = 6 years old